# Dynamics in an emergent quantum-like state space generated by a nonlinear classical network

### v1

@@ Abstract
This work exploits a framework whereby a graph (in the mathematical sense) serves to connect a classical system to a state space that we call ‘quantum-like’ (QL). The QL states comprise arbitrary superpositions of states in a tensor product basis. The graph plays a special dual role by directing design of the classical system and defining the state space. We study a specific example of a large, dynamical classical system—a system of coupled phase oscillators—that maps, via a graph, to the QL state space. We investigate how mixedness of the state diminishes or increases as the underlying classical system synchronizes or de-synchronizes respectively. This shows the interplay between the nonlinear dynamics of the variables of the classical system and the QL state space. We prove that maps from one time point to another in the state space are linear maps. In the limit of a strongly phase-locked classical network—that is, where couplings between phase oscillators are very large—the state space evolves according to unitary dynamics, whereas in the cases of weaker synchronization, the classical variables act as a hidden environment that promotes decoherence of superpositions.
@@ I Introduction
In recent work we have described how a classical system can be mapped, via a graph representation, to a state space that has many properties similar to a quantum state space[1, 2]. For that reason, we call it a quantum-like (QL) state space, and refer to the states as QL states. The graphs needed as a basis fgor these constructions are called QL bits. Particular maps combine these special graphs to produce states that, remarkably, have the properties of quantum states. In particular, by applying a product operation to QL bit graphs we can produce graphs whose states comprise arbitrary (controllable) superpositions of states in a tensor product basis. These same graphs can also be used to define classical networks. In the present paper we study how dynamics of the classical networks are reflected in the properties of the QL states.
@@
We study explicit examples where graphs map to classical states that are networks of coupled phase oscillators, therefore we can use the well-developed theories for synchronization phenomena to provide dynamical models for the classical network. The QL graph construction allows us to add a layer to the dynamical network—an associated state space in the tensor product basis. This is where the new results of this work lie. We explore the way synchronization and de-synchronization of the classical system encodes dynamics in the state space that can be described as purity-preserving dynamics and decoherence, respectively.
@@
The QL bit graphs can represent physical, classical systems, most obviously networks of phase oscillators, as studied here. The crucial feature of QL bit graphs and the corresponding classical networks is that they encode two states. In our model, each state can be the emergent state of a subgraph of the QL bit graph, or we can choose the superpositions of emergent states as a basis. QL states, more generally, are represented by product graphs. Generally, these product graphs can be considered to be abstract systems, but they can also be produced by physical systems, such as networks. Here we make no assumptions about the physical manifestation of the phase oscillator system underlying the QL bits, but we could imagine it to represent some kind of dynamical structure of fields or particles that collectively form a microscopic QL object[3, 4, 5]. Or, more abstractly, these underlying variables might be thought of as an intuitive way of representing wavelike properties of a quantum particle because they capture the nature of phase. On the other hand, it can simply be a macroscopic classical network, for which there are numerous examples[6, 7]. Regardless of these possible interpretations, here we study the specific case of coupled phase oscillators serving as a classical system defied by a QL state graph, and a corresponding state space generated by the adjacency matrix of that graph.
@@
Figure 1: Depiction of how the QL state graphs serve to direct, on one hand, the construction of a representative classical system, and on the other hand, a corresponding QL state space.
@@
In Fig. 1 the correspondence between the classical system of phase oscillators and the graph and the QL state space is shown. The graph plays a special dual role by directing design of the classical system and defining the state space. The QL state space, specifically, comprises the set of states, projected out of the eigenvectors of the adjacency matrix, that are indicated by the emergent states, highlighted by the box over the spectrum. The interplay between the classical system and these states comes about in the the present work by the evolution of the oscillator phases. We say more about how we configure the classical system in relation to the state space in the next section of the paper.
@@
We develop the idea that there are two viewpoints for our system. First, we can observe the nonlinear dynamics of the complex classical network either at a very detailed classical level, or by considering a global order parameter. Second, we can study the state space in the reduced representation of the emergent QL states. The graph model for QL states[1] connects these two viewpoints. What seems particularly interesting about the parallel view of the same dynamics, is that it offers a way to think about the underlying classical variables that control properties of the QL state space—those variables can be hidden if we imagine we can only observe the state space and do not know much, or anything, about the classical system that is controlling the states. However, they are not necessarily deterministic, depending how we define the averaging (see the next section).
@@ II How graphs map classical variables to the state space
To start, we briefly provide some background. A graph $G(n,m)$, or simply $G$, comprises $n$ vertices and a set of $m$ edges that connect pairs of vertices. The size of a graph or subgraph, that is, the number of vertices, is written $|G|$. A graph has two important features. First, the structure of the graph—determined by the pattern of edges—plays a key role in the classical dynamical evolution of a system represented by the graph. That is because the structure of couplings between phase oscillators in the network iplays an important role in determining phase synchronization[8, 9, 10, 11]. Second, any graph can be characterized by its spectrum and associated eigenstates. Each (finite) graph is represented by a matrix called the adjacency matrix, $A$, and the spectrum of a graph is obtained by diagonalizing the corresponding adjacency matrix. The adjacency matrix is indexed by the vertices and (for a normal graph) contains an entry 1 at $a_{ij}$ if the vertex $i$ is linked by en edge to vertex $j$. We can assign a phase (or bias) to the edges so that they may take any value on the unit circle in the complex plane[12, 13, 14]. We will exploit this freedom in the present work to map a classical network of phase oscillators to a graph Here we will specialize to undirected graphs, so the adjacency matrix is symmetric ($a_{ij}=a_{ji}$). For more specialized background on graph theory see [15, 16, 17].
@@
Special graphs, called expander graphs, have a spectrum comprising a distinguished emergent state—a state that is clearly separated from other states in the spectrum. The emergent state of an expander graph is the resource we use to construct the QL state space. Two expander graphs are bound by connecting edges to produce a two-state system that we call the QL bit[1]. These QL bit graphs are combined using a graph operation—the Cartesian product of graphs—to generate a new, exponentially larger, graph. The state states of this product graph are linear combinations of tensor products of the states of the QL bits[2]. For background see ref [18].
Hidden variable theories in quantum mechanics aim to enable a physical interpretation of quantum mechanics, in particular to address questions such as how measurement works, what is the relationship between quantum and classical theories, and so on[19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. The main scope of hidden variable theories is to ask whether there is a deterministic system with predetermined outcomes prescribed for its properties, but this picture is hidden to us by the ways we measure the state vector. The motivation is that the hidden picture would clarify some of the mysteries that underpin the quantum theory. The no-hidden-variables theorems show how this is not the case[20].
Another program is to understand how superpositions might collapse under dynamics governed by a modified version of the Schrödinger equation[33, 34, 35, 3]. Here the challenge is to understand how superpositions are lost, which is impossible under the linear evolution of the Schrödinger equation. Solving that problem would explain why macroscopic systems are not found in superposition and the theory would thus demystify the quantum to classical transition. One proposal is that the Schrödinger equation is modified so that at the micro-scale it appears unchanged from our usual formulation, but at longer length scales a nonlinear, stochastic feature emerges[34]. This new feature acts to collapse superpositions. In this program, details of the nonlinear feature are hidden when we measure the state vectors.
In the present work, we study a different way of imagining a hidden nonlinear system. By connecting a classical system to a state space (here the QL state space), we can explore how concrete properties of the classical system affect the states. We consider networks of phase oscillators for the classical system, so the implications of phase in the state vectors comes to the forefront. We can ask, for example, whether collapse of superposition is possible in an isolated system because of the way the states are generated by a system that cannot lock the phases of the states? To explore this question we set up a map from a complex classical system to a quantum-like state space that is mediated by properties of a suitable graph. An appropriate state space is provided by the graph representation for a QL bit[1], combined with the Cartesian product operation on the graphs to produce a new graph that generates a state space mimicking arbitrary quantum states[2]. The classical system we consider here, as mentioned, is a network of phase oscillators. The structure of that network is templated by the graphs that generate the QL state space.
@@
Any QL state space is mirrored by a corresponding graph. That graph can be considered to be an abstract concept, but can also be physically realized. A notable example is the system of phase oscillators studied here. Each oscillator is specified by a vertex in the graph and the edges show which phase oscillators are pairwise coupled. Notice here that the classical system is related one-to-one with the graph structure, although this is not a general requirement of the mapping of the classical system to the state space via a graph. The graph, or equivalently the network of phase oscillators, is much more complex than the state space because we reduce the state space to include only the emergent states (Sec. IV.B). The many other ‘random’ states in the spectrum[18] are ignored. Thus, the complexity of the graph, or classical system, is reduced in the QL state space.
The evolution of the oscillator system, though sensitive to the specified initial conditions, is entirely deterministic. However, we use a particular averaging over the initial conditions that reflects our hypothesis that we cannot know the initial phases of the oscillators, only the variance of the distribution. Therefore, in the numerical studies we take an ‘ensemble’ average over initial conditions of the phase oscillators which is intended to capture the fact that we cannot know precisely the phases. It is not necessarily an average over a set of distinct QL systems. Details of the classical system are therefore obscured in the state space by the unknown phases of the oscillators.
In this report, we study how dynamics of the classical system compares to, and governs, dynamics in the QL state space. The basis for the study are the nonlinear dynamics of a large, complex classical network of phase oscillators. We ask, how are these dynamics reflected in the reduced space of QL states?
@@ III Synchronization
The literature on synchronization phenomena is vast[6, 36, 37, 38, 39, 40, 41, 42]. The idea of collective motion is well known to us from everyday experiences, and provides a good concrete example of physical synchronization[43, 44]. There are numerous examples in biology, from circadian rhythms to cell cycles to neural networks to ensembles of atoms[45, 46, 47, 48, 49, 50, 51].
@@
In the case of a system of many interacting oscillators, the Kuramoto model[39, 36] attributes the nonlinearity to the coupling among the oscillators:
@@
where $\nu_{i}$ is the frequency of the oscillator at node $i$, $\theta_{i}$ is the oscillator phase defined in terms of accumulated phase and an offset $\theta_{i}(t)=\epsilon_{i}t+\phi_{i}$. $K$ is the coupling value, $a_{ij}$ (with value 0 or 1) are taken from the adjacency matrix of the graph representing the node connectivities in the network, and the coupling is renormalized by the factor $1/N$ so that networks are size-intensive with respect to the lowest eigenvalue.
This remarkably simple system of equations predicts how oscillators collectively synchronize, according to the chosen parameters such as frequency disorder, coupling strength, and initial phases of the oscillators. Feedback introduced by small coupling between pairs of oscillators causes oscillators to speed up or slow down during a period of oscillation—according to their natural frequency relative to the mean—ultimately equilibrating to a steady state where, for some initial conditions, phases remain locked in step. It is entropically favorable to minimize $\theta_{ij}=\theta_{j}-\theta_{i}$ because otherwise the system has to maintain a sequence of correlations among the evolving phases, decided by the sum rule for angles in an $n$-gon, $\sum\theta_{ij}=\pi(n-2)$. That influences the allowed phase differences. For example, consider three oscillators in the network ($i,j,k$), where we know $\theta_{ij}$ and $\theta_{ik}$, then we must have that $\theta_{kj}=\pi-\theta_{ij}-\theta_{ik}$. Thus many correlations are embedded in the coupling matrix.
The propensity for a system of oscillators to synchronize depends on the form of the coupling matrix, that is, the structure of the graph where each oscillator is represented by a vertex and pairwise coupling is indicated by an edge[52, 8]. Synchronization also tends to be more robust when many oscillators are coupled, provided that the couplings are global. Notably, systems with only nearest-neighbor coupling synchronize poorly compared to more highly connected networks[9, 52].
@@
A connection between this synchronization model and random matrix theory was established in ref [53]. In that work we considered a network of phase oscillators defined by a graph and considered a Gaussian distribution of frequencies for the oscillators. It was noted that an ensemble of synchronization trajectories, each started with a random phase distribution for the oscillators, equilibrated to a phase distribution representative of the order parameter and indicative of the standard deviation of the frequency distribution. That is, for example, when the standard deviation of the frequency distribution is less than the coupling $K$, but not much less, then the equilibrium phases are clustered with a certain circular standard deviation, which can be related to the frequency distribution. It was thus proposed that the equilibrium properties of the phase oscillator ensemble can be predicted using a phase-oscillator version of random matrix theory. A relevant finding is that the states of a poorly synchronized phase oscillator network are more mixed than those of a synchronized network. We use a related method in the calculations of spectra and states described below.

### v3

@@ Abstract
This work exploits a framework whereby a graph (in the mathematical sense) serves to connect a classical system to a state space that we call ‘quantum-like’ (QL). The QL states comprise arbitrary superpositions of states in a tensor product basis. The graph plays a special dual role by directing design of the classical system and defining the state space. We study a specific example of a large, dynamical classical system—a system of coupled phase oscillators—that maps, via a graph, to the QL state space. We investigate how mixedness of the state diminishes or increases as the underlying classical system synchronizes or de-synchronizes respectively. This shows the interplay between the nonlinear dynamics of the variables of the classical system and the QL state space. We prove that maps from one time point to another in the state space are linear maps. In the limit of a strongly phase-locked classical network—that is, where couplings between phase oscillators are very large—the state space evolves according to unitary dynamics, whereas in the cases of weaker synchronization, the classical variables act as a hidden environment that promotes decoherence of superpositions. We examine how similar the properties of QL states are to quantum states. We find that during decoherence of the QL states, the off-diagonal density matrix elements decay and that this decay can be observed in any basis we choose for measurement. More surprisingly, we show that a no-cloning theorem (that is, a state of a QL bit cannot be copied) applies not only to the QL states, but also to the underlying classical system.
@@ I Introduction
In recent work we have described how a classical system can be mapped, via a graph representation, to a state space that has many properties similar to a quantum state space[1, 2]. For that reason, we call it a quantum-like (QL) state space, and refer to the states as QL states. The graphs needed as a basis for these constructions are called QL bit graphs (or simply QL bits). By applying a product operation to QL bit graphs we can produce graphs whose states comprise arbitrary (controllable) superpositions of states in a tensor product basis—entangled states. These same graphs can also be used to define classical networks. In the present paper we study how dynamics of the classical networks are reflected in the properties of the QL states.
@@
The main aim of this report is to connect the abstract formalism for QL states we developed in prior work to an example of a physical model represented by the graphs. We study explicit examples where graphs map to classical states that are networks of coupled phase oscillators, therefore we can use the well-developed theories for synchronization phenomena to provide dynamical models for the classical network. The QL graph construction allows us to add a layer to the dynamical network—an associated state space in the tensor product basis. This is where the new results of this work lie.
The physical motivation for the study is to investigate, using numerical studies, how the QL states evolve as a result of evolution of the underlying classical network. We explore the way synchronization and de-synchronization of the classical system translates to dynamics in the state space that can be described as purity-preserving dynamics and decoherence of QL states, respectively. We are motivated to ask how similar the properties of QL states are to quantum states. We find that during decoherence of the QL states, the off-diagonal density matrix elements decay and that this decay can be observed in any basis we choose for measurement. More surprisingly, we show that a no-cloning theorem (that is, a state of a QL bit cannot be copied) applies not only to the QL states, but also to the underlying classical system.
@@
The QL bit graphs can represent physical, classical systems. Most obviously these can be networks of phase oscillators, as studied here. The crucial feature of QL bit graphs and the corresponding classical networks is that they encode two states. In our model, each state can be the emergent state of a subgraph of the QL bit graph, or we can choose the superpositions of emergent states as a basis. QL states, more generally, are represented by product graphs. Here we make no assumptions about the physical manifestation of the phase oscillator system underlying the QL bits, but we could imagine it to represent some kind of dynamical structure of fields or particles that collectively form a microscopic QL object[3, 4, 5]. Or, more abstractly, these underlying variables might be thought of as an intuitive way of representing wavelike properties of a quantum particle because they capture the nature of phase. On the other hand, the phase oscillator system can simply be a macroscopic classical network, for which there are numerous examples[6, 7]. Regardless of these possible interpretations, here we study the specific case of coupled phase oscillators serving as a classical system defined by a QL state graph, and a corresponding state space generated by the adjacency matrix of that graph.
@@
Figure 1: Depiction of how the QL state graphs serve to direct, on one hand, the construction of a representative classical system, and on the other hand, a corresponding QL state space. The $\Box$ symbol denotes the Cartesian product of graphs.
@@
In Fig. 1 the correspondence between the classical system of phase oscillators and the graph and the QL state space is shown. The graph plays a special dual role by directing design of the classical system and defining the state space. The QL state space, specifically, comprises the set of states, projected out of the eigenvectors of the adjacency matrix, that are indicated by the emergent states. These emergent states are highlighted by the box over the spectrum. The interplay between the classical system and these states comes about in the the present work by the evolution of the oscillator phases. We say more about how we configure the classical system in relation to the state space in the next section of the paper.
@@
We develop the idea that there are two viewpoints for our system. First, we can observe the nonlinear dynamics of the complex classical network either at a very detailed classical level, or by considering a global order parameter. Second, we can study the state space in the reduced representation of the emergent QL states. The graph model for QL states[1] connects these two viewpoints.
Two questions need clarification. First, if we represent the QL graph as a network (in the present work, a network of oscillators), what is the difference between the graph and the classical network, because they appear to look the same? The graph encodes rules, according to those rules the classical network is constructed. We can copy the structure of the graph to the network, like we do here. But note that the network needs other features including an assignment of frequencies to the vertices and, possibly, weighting of the edges (couplings). However, we can abstract the design rules and choose a classical system with many more vertices or a continuum of vertices—as long as the structure of each subgraph is $d$-regular then the spectrum will be the same as the graph. Notably, the classical system does not have a state space, but the graph can generate the state space we need via its adjacency matrix.
Second, this state space is defined by the adjacency matrix associated with the graph, so what is the difference between the graph and the QL states? The states are generated by the graph via its adjacency matrix, but they do not uniquely define the graph and its physical properties, as noted below. If we know a gate operation or unitary transformation we want to perform on the QL states, we can find an operation on the QL graph that produces the desired transformation in the state space[8]. However, it is not always obvious how the graph should be transformed. For example, a unitary map on a state does not correspond to an obvious structural change, like an isomophism on the graph. The unitary map operates on the emergent QL states, like those defined in Sec. IV.B, and is defined by a linear map preserving inner product. The graph isomorphism is a bijective map between vertex sets that preserves adjacency and thereby produces two distinct graphs with the same properties. As long as we concomitantly redefine the projections to the QL basis, the states are unchanged.
What seems particularly interesting about the parallel view of the same dynamics, is that it offers a way to think about the underlying classical variables that control properties of the QL state space—those variables can be hidden if we imagine we can only observe the state space and do not know much, or anything, about the classical system that is controlling the states. However, they are not necessarily deterministic, depending how we define the averaging over the system preparation (see the next section).
@@ II How graphs map classical variables to the state space
To start, we briefly provide some background. A graph $G(n,m)$, or simply $G$, comprises $n$ vertices and a set of $m$ edges that connect pairs of vertices. The size of a graph or subgraph, that is, the number of vertices, is written $|G|$. A graph has two important features. First, the structure of the graph—determined by the pattern of edges—plays a key role in the classical dynamical evolution of a system represented by the graph. That is because the structure of couplings between phase oscillators in the network plays an important role in determining phase synchronization[9, 10, 11, 12]. Second, any graph can be characterized by its spectrum and associated eigenstates, but not uniquely[13, 14]. In the present context, for instance, it is possible to find various different (not isomorphic) random graphs, that are on average $d$-regular, with the same spectrum.
Each (finite) graph is represented by a matrix called the adjacency matrix, $A$, and the spectrum of a graph is obtained by diagonalizing the corresponding adjacency matrix. The adjacency matrix is indexed by the vertices and (for a normal graph) contains an entry 1 at $a_{ij}$ if the vertex $i$ is linked by en edge to vertex $j$. We can assign a phase (or bias) to the edges so that they may take any value on the unit circle in the complex plane[15, 16, 17]. We will exploit this freedom in the present work to map a classical network of phase oscillators to a graph Here we will specialize to undirected graphs, so the adjacency matrix is symmetric ($a_{ij}=a^{*}_{ji}$, where * means complex conjugate). For more specialized background on graph theory see [18, 19, 20].
@@
Special graphs, called expander graphs, have a spectrum comprising a distinguished emergent state—a state that is clearly separated from other states in the spectrum. The emergent state of an expander graph is the resource we use to construct the QL state space. Two expander graphs are bound by connecting edges to produce a two-state system that we call the QL bit[1]. These QL bit graphs are combined using a graph operation—the Cartesian product of graphs—to generate a new, exponentially larger, graph. The states of this product graph are linear combinations of tensor products of the states of the QL bits[2]. For background see ref [21].
@@
To explore a concrete realization of a classical system generating QL states, we set up a map from a complex classical system to a quantum-like state space that is mediated by properties of a suitable graph. The classical system we consider here, as mentioned, is a network of phase oscillators. The structure of that network is templated by the graphs that generate the QL state space. The graph, or equivalently the network of phase oscillators, is much more complex than the state space because we reduce the state space to include only the emergent states (Sec. IV.B). The many other ‘random’ states in the spectrum[21] are ignored. Thus, the complexity of the graph, or classical system, is reduced in the QL state space by shifting our focus to the emergent states.
We consider networks of phase oscillators for the classical system, so the implications of phase in the state vectors comes to the forefront. For instance, we show in the results of this work how phase differences lead to decoherence of states when the oscillator network fails to remain synchronized. We thus find that collapse of superposition is possible in an isolated system because of the way the states are generated by a system that cannot lock the phases of the states. We further show how the preferred basis (or pointer states[22]) are defined by the way we observe the QL states.
Any QL state space is mirrored by a corresponding graph in the sense that the graph dictates certain design rules that the classical system must satisfy. For example, we require subsystems that produce emergent states and these subsystems are coupled in a particular way so that correlations are built into the system that match those defining a quantum state space. The classical system should not be confused with an arbitrary classical system. For instance, if the states obey a no-cloning theorem (which they do), then this cannot be circumvented by an allowed transformation of the classical network (see Sec. V.F).
In this report, we study how dynamics of the classical system compares to, and governs, dynamics in the QL state space. The basis for the study are the nonlinear dynamics of a large, complex classical network of phase oscillators. We ask, how are these dynamics reflected in the reduced space of QL states? The work substantiates the claim made in our prior work that QL states can be robust to disorder and resist decoherence, which is demonstrated here by synchronization conditions for the classical network. In contrast, when the classical network does not synchronize, then an initially prepared synchronized system produces QL states that become more mixed with time—they show decoherence. It is shown how we can observe that decoherence in any basis we choose according to our measurement. In the case of QL states undergoing decoherence, we can think of the poorly-synchronized classical network as a kind of intrinsic bath.
@@ III Synchronization
The literature on synchronization phenomena is vast[6, 23, 24, 25, 26, 27, 28, 29]. The idea of collective motion is well known to us from everyday experiences, and provides a good concrete example of physical synchronization[30, 31]. There are numerous examples, from circadian rhythms to cell cycles to neural networks to ensembles of atoms[32, 33, 34, 35, 36, 37, 38].
@@
In the case of a system of many interacting oscillators, the Kuramoto model[26, 23] attributes the nonlinearity to the form of coupling among the oscillators:
@@
where $\nu_{i}$ is the frequency of the oscillator at node $i$, $\theta_{i}$ is the oscillator phase defined in terms of accumulated phase and an offset $\theta_{i}(t)=\epsilon_{i}t+\phi_{i}$. $K$ is the coupling value, $a_{ij}$ (with value 0 or 1) are taken from the adjacency matrix of the graph. The $a_{ij}$ represent the node connectivities in the network. The coupling is renormalized by the factor $1/N$ so that networks are size-intensive with respect to the lowest eigenvalue. Notice that the nonlinearity comes about because of the terms $f(\theta_{j}-\theta_{i})$, where $f(\dots)=\sin(\dots)$.
This remarkably simple system of equations predicts how oscillators collectively synchronize, according to the chosen parameters such as frequency disorder, coupling strength, and initial phases of the oscillators. Feedback introduced by small coupling between pairs of oscillators causes oscillators to speed up or slow down during a period of oscillation—according to their natural frequency relative to the mean. Ultimately, the oscillator system equilibrates to a steady state where, for some initial conditions, phases remain locked in step. It is entropically favorable to minimize $\theta_{ij}=\theta_{j}-\theta_{i}$ because otherwise the system has to maintain a sequence of correlations among the evolving phases, decided by the sum rule for angles in an $n$-gon, $\sum\theta_{ij}=\pi(n-2)$. That influences the allowed phase differences. For example, consider three oscillators in the network ($i,j,k$), where we know $\theta_{ij}$ and $\theta_{ik}$, then we must have that $\theta_{kj}=\pi-\theta_{ij}-\theta_{ik}$. Thus many correlations are embedded in the coupling matrix.
The propensity for a system of oscillators to synchronize depends on the form of the coupling matrix, that is, the structure of the graph where each oscillator is represented by a vertex and pairwise coupling is indicated by an edge[39, 9]. Synchronization also tends to be more robust when many oscillators are coupled, provided that the couplings are global. Notably, systems with only nearest-neighbor coupling synchronize poorly compared to more highly connected networks[10, 39]. We discuss synchronization in the context of choice of parameters in Sec. V.B.
@@
A connection between this synchronization model and random matrix theory was established in ref [40]. In that work we considered a network of phase oscillators defined by a graph and considered a Gaussian distribution of frequencies for the oscillators. It was noted that an ensemble of synchronization trajectories, each started with a random phase distribution for the oscillators, equilibrated to a phase distribution representative of the order parameter and indicative of the standard deviation of the frequency distribution. That is, for example, when the standard deviation of the frequency distribution is less than the coupling $K$, but not much less, then the equilibrated phases are clustered, with a certain circular standard deviation which can be related to the frequency distribution. It was thus proposed that the equilibrium properties of the phase oscillator ensemble can be predicted using a phase-oscillator version of random matrix theory. A relevant finding is that the states of a poorly synchronized phase oscillator network are more mixed than those of a synchronized network. We use a related method in the calculations of spectra and states described below.
We can also consider systems comprising an infinite number of oscillators. A continuum limit of the Kuramoto model can be developed in terms of oscillator phase densities[25, 41] and Sec. 7 of Ref. [23].